# Video Understanding with Large Language Models: A Survey

### v4

@@ III-B2 Fine-tuning Vid-LLMs
In contrast to most Vid-LLMs in the Video Analyzer $\times$ LLM category being training-free, almost all Vid-LLMs in the Video Embedder $\times$ LLM category undergo fine-tuning. The common methods for fine-tuning Vid-LLMs are categorized based on the types of adapters used during fine-tuning into four main types: LLM Fully Fine-tuning, Connective Adapter Fine-tuning, Insertive Adapter Fine-tuning, and Fine-tuning with Hybrid Adapters. An adapter is a small, trainable module added to a large model for fine-tuning. By only updating the parameters of these modules, the model can adapt to specific tasks or domains without changing the entire model’s parameters, achieving efficient parameter updates and task adaptation while conserving computational and storage resources. Illustrations of each type are shown in Figure 6.
@@
  • Fine-tuning with Hybrid Adapters: Many Vid-LLMs use a combination of Connective and Insertive Adapters to achieve both modality alignment and changes in the LLM’s inherent behavior. Vid-LLMs employing Hybrid Adapters typically use multi-stage fine-tuning. A common approach is to fine-tune only the Connective Adapter in the first stage for modality alignment. In the second stage, the already fine-tuned Connective Adapter is frozen, the training task (from alignment task to target task) and the training data (from data used for modality alignment to data required for the target task) are changed, and only the parameters of the Insertive Adapter are updated. There are also single-stage approaches where both Connective Adapters’ and Insertive Adapters’ parameters are updated simultaneously.
@@ IV Benchmarks and Evaluation
This section provides an overview of the evaluation methods for video question-answering models and related tasks, categorized into three types: closed-set evaluation, open-set evaluation, and other evaluation methods. Closed-set evaluation relies on questions with predefined answers, measuring model performance by calculating the percentage of correct answers. It uses metrics like CIDEr [279], METEOR [280], ROUGE [281] and SPICE [282] to compare predicted results with ground truth. Open-set evaluation does not rely on predefined options but still requires ground truth answers for scoring, often using models like GPT-3.5/4 to compare predictions and assign scores. Other evaluation methods include tasks requiring fine-grained temporal and spatiotemporal understanding, such as dense captioning, video temporal grounding, spatiotemporal grounding, and object tracking, using metrics like IoU, Recall@K, and mAP. Additionally, human evaluation remains a common approach, though labor-intensive and time-consuming.
IV-A Closed-set Evaluation
Closed-set evaluations are based on questions with pre-defined answers [283]. For video question-answering tasks, the questions are usually designed as multiple-choice, and the performance of models can be easily evaluated by calculating the percentage of correct answers, resulting in an accuracy score. For captioning or video-to-text summarization tasks, the ground truth of the captions or summaries is pre-prepared. The CIDEr [279], METEOR [280], ROUGE [281] and SPICE [282] metrics are computed by comparing the predicted sentences with the ground-truth sentences. The relevant benchmarks include MSRVTT-QA [93], MSVD-QA [93], TGIF-QA [107], ActivityNet-QA [110], TVQA [109], How2QA [284], STAR [285], NExT-QA [111], MVBench [197], Video-Bench [286], EgoSchema [287], AutoEval-Video [288], TempCompass [289], Video-MME [290], VideoVista [291], CinePile [292], SOK-Bench [293], SFD [294], EditVid-QA [295], InfiniBench [296], MLVU [297], and MMWorld [298]. The detailed information of these benchmarks is shown in Table III.
@@ IV-A Closed-set Evaluation
TABLE III: The comparison of various benchmarks includes several important aspects: the total number of videos, the number of clips, the average duration of the videos, and the number of QA pairs.
Benchmarks #Videos #Clips Len.(s) #QA Pairs
MSRVTT-QA [93] 2,990 2,990 15.2 72,821
MSVD-QA [93] 504 504 9.8 13,157
TGIF-QA [107] 9,575 9,575 3.0 8,506
ActivityNet-QA [110] 800 800 111.4 8,000
TVQA [109] 2,179 15,253 11.2 15,253
How2QA [284] 1,166 2,852 15.3 2,852
STAR [285] 914 7,098 11.9 7,098
NExT-QA [111] 1,000 1,000 39.5 8,564
MVBench [197] 3,641 3,641 16.0 4,000
Video-Bench [286] 5,917 5,917 56.0 17,036
EgoSchema [287] 5,063 5,063 180.0 5,063
AutoEval-Video [288] 327 327 14.6 327
TempCompass [289] 410 500 11.4 7,540
Video-MME [290] 900 900 1,017.9 2,700
VideoVista [291] 894 3,400 131.0 25,000
CinePile [292] 9,396 9,396 160 303,828
SOK-Bench [293] 10,000 10,000 - 44,000
SFD [294] 1,078 1,078 780 4,885
EditVid-QA [295] 32,000 32,000 - 204,000
InfiniBench [296] 1,219 1,219 4,460.4 108,200
MLVU [297] 1,334 1,334 180-7,200 2,593
MMWorld [298] 1,910 1,910 102.3 1,599
VELOCITI [299] 900 900 10.0 -
IV-B Open-set Evaluation
Unlike closed-set evaluation, open-set evaluation does not rely on pre-defined options. Ground-truth answers are still needed, but the scores are not derived from counting correct options or calculating the similarity between the predicted and ground-truth sentences. Instead, models like GPT-3.5/4 are prompted to compare the models’ predictions with the answers and assign scores to the predictions. Some commonly used datasets or benchmarks for video question-answering are MSRVTT-QA [93], MSVD-QA [93], TGIF-QA [107], and ActivityNet-QA [110]. The most popular GPT-based open-set evaluation methods are proposed in [99]: Open-end Zero-shot Video QA Evaluation and Video-based Generative Performance Benchmarking. We collect the performance of some popular Vid-LLMs on these evaluation metrics for comparison, as shown in Table IV.
IV-C Others
There are also other benchmarks and evaluation metrics. For instance, benchmarks for tasks that assess models’ fine-grained temporal and spatiotemporal understanding capacities, such as dense captioning, video temporal grounding, spatiotemporal grounding, object tracking, video saliency detection, etc., require temporal or spatiotemporal annotations as ground-truth to evaluate performance, using metrics like IoU, Recall@K, and mAP. Human evaluation is also a commonly used method for evaluation, although it is labor-intensive and time-consuming.
@@ IV-C Others
TABLE IV: This table comprehensively compares various Vid-LLMs across multiple open-end zero-shot video question answering and video-based generative performance benchmarks. It includes GPT-based metrics for MSVD-QA, MSRVTT-QA, and ActivityNet-QA datasets, as well as scores for correctness, detail, context, temporal, and consistency aspects in video-based generative performance.
@@
Model MSVD-QA MSRVTT-QA ActivityNet-QA Correctness Detail Context Temporal Consistency
GPT4-V - - 59.5 4.09 3.88 4.37 3.94 4.02
Video-LLaMA 51.6 29.6 12.4 1.96 2.18 2.16 1.82 1.79
LLaMA-Adapter 54.9 43.8 34.2 2.03 2.32 2.30 1.98 2.15
VideoChat 56.3 45.0 26.5 2.23 2.50 2.53 1.94 2.24
VALLY 60.5 51.1 45.1 2.43 2.13 2.86 2.04 2.45
MovieLLM 63.2 52.1 43.3 2.64 2.61 2.92 2.03 2.43
Video-ChatGPT 64.9 49.3 35.2 2.50 2.57 2.69 2.16 2.20
Chat-UniVi 65.0 54.6 46.1 2.89 2.91 3.46 2.89 2.81
Vista-LLaMA 65.3 60.5 48.3 2.44 2.64 3.18 2.26 2.31
AV-LLM 67.3 53.7 47.2 2.56 2.47 2.93 2.17 2.47
LLaVA-NeXT-Video 67.8 - 53.5 3.39 3.29 3.92 2.60 3.12
LLaMA-VID 69.7 57.7 47.4 2.96 3.00 3.53 2.46 2.51
VTimeLLM 69.8 58.8 45.5 2.78 3.10 3.40 2.49 2.47
VideoChat2 70.0 54.1 49.1 3.02 2.88 3.51 2.66 2.81
LongVLM 70.0 59.8 47.6 2.76 2.86 3.34 2.39 3.11
AVicuna 70.2 59.7 53.0 2.81 2.62 3.25 2.53 2.59
Video-LLaVA 70.7 59.2 45.3 2.87 2.94 3.44 2.45 2.51
RED-VILLM 71.2 53.9 44.2 2.71 2.75 3.34 2.34 2.45
Video LLaMA 2 71.7 - 49.9 3.09 3.09 3.68 2.63 3.25
Artemis 72.1 56.7 39.3 2.69 2.55 3.04 2.24 2.70
VideoGPT+ 72.4 60.6 50.6 2.46 2.73 2.81 1.78 2.47
MiniGPT4-video 73.9 58.8 44.4 3.09 3.02 3.57 2.56 2.67
ST-LLM 74.6 63.2 50.9 3.23 3.05 3.74 2.93 2.81
MovieChat 75.2 52.7 45.7 2.76 2.93 3.01 2.24 2.42
PLLaVA 76.6 62.0 56.3 3.21 2.86 3.62 2.33 2.93
IG-VLM 76.7 62.7 57.3 3.26 2.76 3.57 2.34 3.28
V Applications
@@
Vid-LLMs have revolutionized various sectors by enabling advanced video and language processing capabilities. This section outlines their diverse applications, demonstrating the extensive and transformative impact of Vid-LLMs across industries.
V-A Media and Entertainment
@@
  • Online Video Platforms and Multimedia Information Retrieval: Vid-LLMs significantly enhance search algorithms [300], generate context-aware video recommendations [301], and aid in natural language tasks such as subtitle generation and translation [266], contributing to online video platforms and multimedia information retrieval systems. Their capabilities in analyzing videos for specific keyword retrieval [162, 302, 303] improve intelligent recommendation systems. In the multimedia fields, it combines videos in domains like music [304], avatar [305, 306, 307, 308], and scene [309], to assist with content generation.
@@
V-B Interactive and User-Centric Technologies

### v5

@@ III-B2 Fine-tuning Vid-LLMs
In contrast to most Vid-LLMs in the Video Analyzer $\times$ LLM category being training-free, almost all Vid-LLMs in the Video Embedder $\times$ LLM category undergo fine-tuning. The common methods for fine-tuning Vid-LLMs are categorized based on the types of adapters used during fine-tuning into four main types: LLM Fully Fine-tuning, Connective Adapter Fine-tuning, Insertive Adapter Fine-tuning, and Fine-tuning with Hybrid Adapters. An adapter is a small, trainable module added to a large model for fine-tuning. By only updating the parameters of these modules, the model can adapt to specific tasks without changing the entire model’s parameters, achieving efficient parameter updates and task adaptation while conserving computational resources. Illustrations of each type are shown in Figure 6.
@@
  • Fine-tuning with Hybrid Adapters: Many Vid-LLMs use a combination of Connective and Insertive Adapters to achieve both modality alignment and changes in the LLM’s inherent behavior. Vid-LLMs employing Hybrid Adapters typically use multi-stage fine-tuning. A common approach is to fine-tune only the Connective Adapter in the first stage for modality alignment. In the second stage, the already fine-tuned Connective Adapter is frozen, the training task (from alignment task to target task) and the training data (from data used for modality alignment to data required for the target task) are changed, and only the parameters of the Insertive Adapter are updated. There are also single-stage approaches where both Connective and Insertive Adapters are updated simultaneously.
@@ IV Benchmarks and Evaluation
This section provides an overview of the evaluation methods for video question-answering models and related tasks, categorized into three types: closed-ended evaluation, open-ended evaluation, and other evaluation methods, which are shown in Table III. Closed-ended evaluations rely on questions with predefined answers or formats, including multiple-choice questions and structured formats that allow for straightforward scoring. Open-ended evaluations involve questions without predefined answer options, often requiring more sophisticated scoring methods, including LLM-based evaluations. Other evaluation methods address specialized video understanding capabilities such as temporal/spatiotemporal reasoning.
IV-A Closed-ended Evaluation
Closed-ended evaluations use pre-defined answers or structured formats [283]. These include multiple-choice questions (TVQA [109], How2QA [284], STAR [285]) and questions with structured formats for direct comparison with ground truth (MSRVTT-QA [93], MSVD-QA [93], MVBench [203]). Multiple-choice performance is evaluated via accuracy percentages, while structured formats use metrics like CIDEr [286], METEOR [287], ROUGE [288], and SPICE [289] to compare predictions with ground truth. Notable benchmarks include MSRVTT-QA [93], TVQA [109], MVBench [203], EgoSchema [290], and Video-MME [291]. Each targets different video understanding aspects: TGIF-QA [107] focuses on action recognition, state transition, frame-level QA, and counting; ActivityNet-QA [110] covers motion, spatial, temporal, and free-form dimensions; VidComposition [292] emphasizes compositional reasoning; while NExT-QA [111] and MLVU [293] include causal and temporal action reasoning. These diverse question types test various reasoning abilities, though many benchmarks still exhibit domain biases toward common scenarios and lack diversity in rare events or unusual contexts.
@@ IV-A Closed-ended Evaluation
TABLE III: The comparison of various benchmarks includes several important aspects: the total number of videos, the number of clips, the average duration of the videos, the number of QA pairs, and video content.
Benchmark #Videos #Clips Len.(s) Video Content #QA Pairs Question Type
MSRVTT-QA [93] 2,990 2,990 15.2 Open-domain 72,821 Closed-ended & open-ended (what/who/how/when/where)
MSVD-QA [93] 504 504 9.8 Open-domain 50,505 Closed-ended & open-ended (what/who/how/when/where)
TGIF-QA [107] 9,575 9,575 3.0 Open-domain 8,506 Closed-ended & open-ended, action, transition, counting
ActivityNet-QA [110] 800 800 111.4 Human activity 8,000 Closed-ended & open-ended (what/who/how/when/where/why)
TVQA [109] 2,179 15,253 11.2 TV show 15,253 Closed-ended, multiple-choice
How2QA [284] 1,166 2,852 15.3 TV episode 2,852 Multiple-choice
STAR [285] 914 7,098 11.9 Human action 7,098 Multiple-choice
NExT-QA [111] 1,000 1,000 39.5 Daily life 8,564 Multiple-choice and open-ended (causal, temporal, descriptive)
MVBench [203] 3,641 3,641 16.0 Open-domain 4,000 Closed-ended (various tasks)
Video-Bench [294] 5,917 5,917 56.0 Open-domain 17,036 Open-ended (various tasks)
EgoSchema [290] 5,063 5,063 180.0 Egocentric activity 5,063 Closed-ended, procedural understanding
AutoEval-Video [295] 327 327 14.6 Open-domain 327 Open-ended evaluation
TempCompass [296] 410 500 11.4 Open-domain 7,540 Both closed-ended & open-ended, temporal reasoning
Video-MME [291] 900 900 1,017.9 Open-domain 2,700 Closed-ended evaluation
VideoVista [297] 894 3,400 131.0 Open-domain 25,000 Open-ended (descriptive, causal, predictive)
CinePile [298] 9,396 9,396 160.0 Movie 303,828 Movie understanding, open-ended
SOK-Bench [299] 10,000 10,000 - Open-domain 44,000 Open-ended, subject-oriented knowledge
SFD [300] 1,078 1,078 780.0 Movies 4,885 Multiple-choice, open-ended
EditVid-QA [301] 32,000 32,000 - Entertainment 204,000 Open-ended, editing techniques
InfiniBench [302] 1,219 1,219 4,460.4 Movie/TV show 108,200 Closed-ended & open-ended, long-form understanding
MLVU [293] 1,334 1,334 180.0-7,200.0 Open-domain 2,593 Both closed-ended and open-ended, long-form understanding
MMWorld [303] 1,910 1,910 102.3 Open-domain 1,599 Open-ended, multimodal evaluation
VELOCITI [304] 900 900 10.0 Movie - Open-ended, visual understanding
VidComposition [292] 982 982 - Movie/TV show 1,706 Multiple-choice, compositional reasoning
EAGLE [305] 36,000 36,000 16.0-76.0 Egocentric activity 400,000 Open-ended, procedural understanding
IV-B Open-ended Evaluation
Open-ended evaluation involves questions without pre-defined options or structured formats. While ground-truth answers serve as references, scoring methods are more sophisticated than option selection or string matching. GPT-3.5/4 models often evaluate predictions by comparing them with reference answers. Notable open-ended benchmarks include MovieChat-1K [241], MLVU [293], NExT-QA [111], VELOCITI [304], and EAGLE [305]. These require more complex responses demonstrating deeper reasoning. CinePile [298] incorporates analytical tasks like character dynamics and narrative analysis. The most popular GPT-based evaluation methods, proposed in [99], are Open-end Zero-shot Video QA Evaluation and Video-based Generative Performance Benchmarking. Performance comparisons of Vid-LLMs on these metrics are shown in Table V. Originally closed-ended benchmarks like MSRVTT-QA [93], MSVD-QA [93], TGIF-QA [107], and ActivityNet-QA [110] can be repurposed as open-ended in GPT-based evaluations, as LLMs generate free-form responses that GPT models compare to references. These methods have limitations: evaluation scores change with GPT version updates, making cross-study comparisons difficult; results depend heavily on prompt engineering; and LLM evaluators may favor responses similar to their generation patterns rather than objectively assessing quality.
IV-C Other Evaluations
Other benchmarks evaluate fine-grained temporal and spatiotemporal understanding. Dense captioning generates [122, 123, 124, 125, 126] detailed descriptions for multiple video events/objects, using BLEU, METEOR, and CIDEr metrics that assess both temporal localization and descriptive accuracy. Vid-LLMs’ performance of dense video captioning on ActivityNet Captions [127] is shown in Table IV. Several Vid-LLMs have already achieved performance comparable to traditional task-specific models in dense video captioning. Video temporal grounding localizes specific moments based on textual queries, evaluated using tIoU and Recall@K. Spatiotemporal grounding extends this to localize in both space and time, assessed via spatiotemporal IoU and mAP. Object tracking [251, 256] follows objects across frames, evaluated using precision, success rate, and tracking accuracy. Video saliency detection [306] identifies visually salient regions, evaluated with AUC-J, NSS, etc. These tasks rely on temporal or spatiotemporal annotations as ground-truth, with metrics like IoU, Recall@K, and mAP widely adopted. Human evaluation is also used for subjective aspects, though this is labor-intensive and time-consuming.
As for qualitative evaluation, several approaches can effectively assess Vid-LLMs’ performance in addition to numerical metrics. Error analysis [203, 307] for open-ended QA and difference comparisons [258, 120] between model outputs and ground truth annotations (e.g., intervals) for temporal/spatiotemporal understanding provide insights into model limitations. Attention visualization [308] reveals what visual elements the models prioritize when generating responses. Self-explanation [307, 292], where models justify their answers for closed-ended benchmarks, offers valuable insights into reasoning processes and potential misconceptions. Human studies, though resource-intensive, remain helpful in finding models that reflect human preferences.
TABLE IV: Comparison of Vid-LLMs and conventional models (non-LLM-based) on dense video captioning models on ActivityNet Captions dataset.
Model CIDEr SODAc METEOR F1
Non-LLM-based
MT [309] 9.3 5.0
PDVC [123] 29.3 6.0 7.6
CM2 [310] 33.1 6.2 8.5 54.2
LLM-based
Momentor [262] 14.9 2.3 4.7
TimeChat [257] 19.0 4.7 5.7 36.9
VTG-LLM [259] 20.7 5.1 5.9 34.8
AVicuna [120] 22.5 5.1 6.5 45.2
TRACE [311] 25.9 6.0 6.4 39.3
VTimeLLM [312] 27.6 5.8 6.8
TRACE-uni [311] 29.2 6.4 6.9 40.4
GIT [313] 29.8 5.7 7.8 50.6
Vid2Seq [271] 30.2 5.9 8.5 51.8
Streaming Vid2Seq [314] 37.8 6.2 10.0 52.9
Streaming GIT [314] 41.2 6.6 9.0 50.9
@@ IV-C Other Evaluations
TABLE V: This table comprehensively compares various Vid-LLMs across multiple open-end zero-shot video question answering and video-based generative performance benchmarks. It includes GPT-based metrics for MSVD-QA, MSRVTT-QA, and ActivityNet-QA datasets, as well as scores for Correctness of Information, Detail Orientation, Contextual Understanding, Temporal Understanding, and Consistency aspects in video-based generative performance.
@@
Model MSVD-QA MSRVTT-QA ActivityNet-QA Correctness Detail Context Temporal Consistency Average
GPT4-V [315] - - 59.5 4.09 3.88 4.37 3.94 4.02 4.06
Video-LLaMA [245] 51.6 29.6 12.4 1.96 2.18 2.16 1.82 1.79 1.98
LLaMA-Adapter [316] 54.9 43.8 34.2 2.03 2.32 2.30 1.98 2.15 2.16
VideoChat [270] 56.3 45.0 26.5 2.23 2.50 2.53 1.94 2.24 2.29
VALLY [244] 60.5 51.1 45.1 2.43 2.13 2.86 2.04 2.45 2.38
MovieLLM [223] 63.2 52.1 43.3 2.64 2.61 2.92 2.03 2.43 2.53
Video-ChatGPT [99] 64.9 49.3 35.2 2.50 2.57 2.69 2.16 2.20 2.42
Chat-UniVi [237] 65.0 54.6 46.1 2.89 2.91 3.46 2.89 2.81 2.99
Vista-LLaMA [236] 65.3 60.5 48.3 2.44 2.64 3.18 2.26 2.31 2.57
AV-LLM [231] 67.3 53.7 47.2 2.56 2.47 2.93 2.17 2.47 2.52
LLaVA-NeXT-Video [317] 67.8 - 53.5 3.39 3.29 3.92 2.60 3.12 3.26
LLaMA-VID [238] 69.7 57.7 47.4 2.96 3.00 3.53 2.46 2.51 2.89
VTimeLLM [258] 69.8 58.8 45.5 2.78 3.10 3.40 2.49 2.47 2.85
VideoChat2 [203] 70.0 54.1 49.1 3.02 2.88 3.51 2.66 2.81 2.98
LongVLM [213] 70.0 59.8 47.6 2.76 2.86 3.34 2.39 3.11 2.89
AVicuna [120] 70.2 59.7 53.0 2.81 2.62 3.25 2.53 2.59 2.76
Video-LLaVA [239] 70.7 59.2 45.3 2.87 2.94 3.44 2.45 2.51 2.84
RED-VILLM [211] 71.2 53.9 44.2 2.71 2.75 3.34 2.34 2.45 2.72
Video LLaMA 2 [201] 71.7 - 49.9 3.09 3.09 3.68 2.63 3.25 3.15
Artemis [191] 72.1 56.7 39.3 2.69 2.55 3.04 2.24 2.70 2.64
VideoGPT+ [200] 72.4 60.6 50.6 2.46 2.73 2.81 1.78 2.47 2.45
MiniGPT4-video [215] 73.9 58.8 44.4 3.09 3.02 3.57 2.56 2.67 2.98
ST-LLM [218] 74.6 63.2 50.9 3.23 3.05 3.74 2.93 2.81 3.15
MovieChat [241] 75.2 52.7 45.7 2.76 2.93 3.01 2.24 2.42 2.67
PLLaVA [217] 76.6 62.0 56.3 3.21 2.86 3.62 2.33 2.93 2.99
IG-VLM [159] 76.7 62.7 57.3 3.26 2.76 3.57 2.34 3.28 3.04
IV-D Analysis of Model Performance
Analyzing the correlation between model attributes and benchmark performance reveals several key factors driving recent improvements in Vid-LLMs. From Tables IV and V, we observe that models built on larger and more recent foundation LLMs (e.g., IG-VLM with 34B parameters) consistently outperform their smaller counterparts, particularly in zero-shot VideoQA tasks. Models employing more powerful visual embedders such as EVA-CLIP or ViT-G architectures (notably in PLLaVA, IG-VLM, and Video LLaMA 2) demonstrate superior performance across both dense captioning and QA benchmarks. The frame sampling strategy also significantly impacts results, with high performers on temporal tasks (like VTimeLLM, AVicuna, and ST-LLM) typically processing more frames (100+) than general understanding models, while sophisticated adaptation mechanisms beyond simple projection layers (such as Q-formers or cross-attention) contribute to better contextual understanding. Performance gains stem from a combination of stronger foundation models, better visual encoders, appropriate temporal modeling, and more sophisticated bridging architectures rather than any single innovation.
V Applications and Future Directions
V-A Application Scenarios
@@
Vid-LLMs have revolutionized various sectors by enabling advanced video and language processing capabilities. We outlines their diverse applications, demonstrating the extensive and transformative impact of Vid-LLMs across industries.
V-A1 Media and Entertainment
@@
  • Online Video Platforms and Multimedia Information Retrieval: Vid-LLMs significantly enhance search algorithms [318], generate context-aware video recommendations [319], and aid in natural language tasks such as subtitle generation and translation [271], contributing to online video platforms and multimedia retrieval systems. Their capabilities in analyzing videos for specific keyword retrieval [168, 320, 321] improve intelligent recommendation systems. In the multimedia fields, it combines videos in domains like music [322], avatar [323, 324, 325, 326], and scene [327], to assist with content generation.
@@
V-A2 Interactive and User-Centric Systems